OpenStudy (anonymous):

Paolo wrote the following equation for the perimeter of a rectangle. P=2(l+w) Which equation is equivalent to the equation Paolo wrote?

OpenStudy (plainntall):

The perimeter is just the total of the sides. Since it is a rectangle the two length sides are the same and the two widths for P=l+l+w+w=2l+2w=2(l+w) are any of those equations in your list?

OpenStudy (anonymous):

yeah, nevermind i have the answer i needed